# Physics-informed machine learning for combustion: A review

### v2

@@ 5.5 Combustion super-resolution via discrete PINNs
As mentioned in Section 2.3.1, discrete PINNs can be applied to SR reconstruction tasks. This field-to-field modeling paradigm is particularly suited for SR reconstruction tasks, offering advantages such as enhanced scalability for high-dimensional data, improved generalization to complex working conditions, and efficient handling of structured grid-based simulation results. Given the inherent complexity of combustion modeling, researchers have further augmented the regular discrete PINNs with generative modeling techniques, such as GANs. Compared to discriminative modeling approaches, generative modeling demonstrates superior capability in capturing small-scale turbulence dynamics 316 and refining predictions of discontinuous features 318. Research on the application of discrete PINNs to combustion problems remains relatively limited. This subsection categorizes the discussion into two parts: subsonic combustion and supersonic combustion.
@@
For subsonic combustion, Bode et al. 316 first used the physics-informed enhanced super-resolution generative adversarial network (PI-ESR-GAN) to model turbulent combustion. As shown in Figure 28, the PI-ESR-GAN framework combines unsupervised deep learning with physics-informed loss functions, employing a two-step training methodology to enhance network generalization, particularly for extrapolation tasks. In decaying turbulence tests involving turbulent mixing, the model demonstrated strong a priori and a posteriori performance, with the physics-informed continuity loss term proving essential. The PI-ESR-GAN methodology has been extended to diverse combustion configurations, demonstrating remarkable versatility. In spray combustion applications, the model was successfully implemented for different cases, with validation against experimental data revealing superior predictive capability compared to classical LES approaches 333. For lean premixed combustion and gas turbine combustors, by developing a modified architecture for finite-rate chemistry flow, the enhanced model demonstrated superior accuracy in both a priori and a posteriori validations while maintaining computational efficiency 334. For turbulent non-premixed flames, a multi-mesh strategy was developed to optimize prediction accuracy across different grid resolutions, enabling reliable extrapolation to untrained Reynolds numbers 335. The approach was subsequently extended to premixed combustion 317. For that, the physical information in the loss function were adjusted, the training process was smoothed, and especially the effects of density changes were considered. When applied to DNS data of turbulent premixed flame kernels, the enhanced model achieved remarkable accuracy in reproducing fine-scale dynamics on coarse computational meshes. This advancement enables more efficient statistical analysis, while significantly reducing computational costs compared to traditional approaches. Using high-fidelity DNS data, Chung et al. 336 benchmarked the scaling behaviour of five DL models on the 3D combustion SR task. One of the models incorporated a physics-based loss (i.e., physics-informed loss in a broad sense), which penalized the gradients of channel variables in a supervised manner. It was demonstrated that the benefits of this loss can persist with increasing model size.
Figure 28: Diagram of the PI-ESR-GAN architecture. Reprinted from Reference 316.
@@
For supersonic combustion, Wang et al. 318 developed the PI-GAN for flow field super-resolution reconstruction on RDCs. The model utilized a GAN architecture enhanced with additional physical loss terms to ensure the outputs adhere to fundamental physical principles. Compared to interpolation and other methods, PI-GAN greatly improved the resolution of the flow field while accurately capturing detailed flow structures. Subsequently, Wang et al. 319 proposed an advanced physics-informed recurrent GAN (PI-RGAN), as shown in Figure 29. This innovative framework incorporated temporal derivatives as intermediate variables during the data generation process within a recurrent architecture, while simultaneously employing spatial derivatives of the output to establish comprehensive physical constraints. Experimental analysis revealed some features of the proposed model. The PI-RGAN model could well capture the detailed flow field structures and temporal characteristics in RDCs. Moreover, the introduction of a recurrent structure in the PI-RGAN model enhanced the model robustness against noisy inputs.
Figure 29: The PI-RGAN for the SR reconstruction of flow fields in RDCs. Reprinted from Reference 319, with permission.
6 PIML for other combustion problems
@@
Some studies on PIML for combustion cannot be classified to the previous two sections but are also important and enlightening. Therefore, they will be introduced in this section, which are categorized into four parts: fire safety, thermoacoustics, combustion diagnostics, and others.
@@ 6.1 Fire safety
Bottero et al. 337 replaced the PDE solver in the fire simulator to achieve real-time prediction of wildfire spread, where the level set equation was utilized. Tests on an ideal case and a real wildfire both demonstrated that the PINNs were able to capture the shape of the firelines. Dabrowski et al. 338 used PINNs to solve the level set equation for modeling the wildfire fire-front. A forecast likelihood was introduced to deal with the issue that PINNs failed to maintain temporal continuity when there are extreme changes in exogenous forcing variables. Moreover, the B-PINN was incorporated to provide UQ in fire-front predictions. The models were successfully applied to modeling two recorded grassland fires. Vogiatzoglou et al. 339 applied PINNs to infer the parameters in the wildfire spread model, consisting of energy and mass fraction equations. The parameters included the dispersion coefficient vector, mean gaseous velocity vector, and the overall heat transfer coefficient, so the number of parameters was three in 1D cases and five in 2D cases. Five cases were tested, including both clean and noisy data, and both synthetic and real-world data, demonstrating the accuracy of parameter inference and wildfire prediction. Some other wildfire-related applications do not focus on the spread prediction. For post-fire applications, Seydi et al. 340 used PIML to estimate the soil burn severity and reveal the intricate relationships between soil and vegetation burn severities. For firefighting, the AI Firefighter 341 was proposed to predict the optimal firefighting strategy given an arbitrary terrain and an arbitrary fire distribution, where the NN combined PIML with a decision-making process.
@@
Other applications are not related to wildfire. Kim et al. 342 used PINNs to model and predict the lithium-ion battery thermal runaway, which is an important topic of fire safety. The PDEs of temperature and lithium concentration served as the governing equations, which was a DR system. The model applicability to semi-supervised learning, unsupervised learning, and surrogate modeling was investigated, showing that PINNs exhibited better performance than purely data-driven NNs. Lu et al. 111 used the PI-DeepONet for surrogate modeling of the radiative heat transfer, where the radiation transfer equation (RTE) and its BCs were considered as the physical losses. The model took the location and propagation direction as trunk net inputs, the parameters (such as the absorption coefficients) as branch net inputs, so the solution operator of the RTE could be approximated. The surrogate model offers significant speedup over traditional methods, thus possessing the potential to replace the traditional RTE solvers in industry-scale fire simulations, where the solution of RTE is often a major performance bottleneck. Parsa et al. 343 applied PINNs to predict the transient burning rates of non-charring materials, where the energy equations in the temperature rise phase and burning phase were considered in the loss function. Results showed that PINNs could effectively capture the intricate dynamics of transient burning behavior and were more efficient than traditional numerical methods.
@@ 6.2 Thermoacoustics
Ozan et al. 344 used PINNs to learn the nonlinear limit cycles and adjoint limit cycles occurred in thermoacoustic oscillations. The physics was informed in two ways: (1) the governing equations that are derived from mass, momentum and energy conservation were integrated into the loss function, and (2) the periodicity was enforced using periodic activation functions in the NN. It was demonstrated that the PDE losses could promote the model performance in noisy-data and scarce-data situations, and only the NN with periodic activation functions could accurately extrapolate to the validation set. Mariappan et al. 345 used PINNs to learn thermoacoustic interactions in a bluff body anchored flame combustor, where three PDEs arising from the acoustic model and the van der Pol (VDP) oscillator model served as the physical losses. Based on the available data of the acoustic pressure fluctuations at three locations and the total flame heat release rate, the model aimed to identify four physical parameters and recover the entire fields of the acoustic pressure, velocity, and total heat release. Five cases were tested in the order of increasing complexity with the final case using the experimental data. The satisfactory results indicated the potential of PINNs to improve existing or design new thermoacoustically stable and structurally efficient combustors. Xie et al. 346 extended the VDP equation by introducing three parameters to enhance the model applicability, and then a PINN was applied to infer these parameters, where the data were generated for a horizontal Rijke tube with acoustically compact heat sources. The optimized PINN-EVDP model was shown to be capable of replicating the thermoacoustic system instability behaviour. Further investigations were conducted to compare the Hopf bifurcation and amplitude death characteristics of the coupled extended VDP system with the coupled theoretical Rijke tube systems, showing the physical significance of the extended VDP system in the field of thermoacoustics.
@@ 6.3 Combustion diagnostics
Wang et al. 347 used PINNs to reconstruct the 3D soot volume fraction in flames, where the NN took the 2D images of the accumulated projection value of the extinction coefficients as input, and output the local 3D extinction coefficients, as shown in Figure 30. The losses included the physical loss of the Lambert-Beer law, the boundary loss that constrained the positivity of the extinction coefficients, and the smoothness loss as a regularization term. Tested on two numerical cases and an experimental case, the model showed superior performance compared to conventional reconstruction methods. Later, Wang et al. 348 used PINNs to simultaneously reconstruct the soot temperature and volume fraction fields in experimental laminar ethylene flames. The inputs of the NN were the 2D matrix of the experimental soot radiation field and flame thickness in the line-of-sight (LoS), and the outputs were soot temperature, volume fraction fields, and an auxiliary variable. The physical loss was a first-order PDE for the auxiliary variable, which was derived from the LoS soot radiation integral equation. It was demonstrated that the physical constraint improved not only the predictive capability but also the adaptability and reliability across various combustion conditions.
Figure 30: Schematic of the Lambert-Beer law PINN (LB-PINN) structure. Reprinted from Reference 347, with permission.
Wang et al. 349 proposed a physics-informed shadowgraph network to reconstruct both density and temperature fields directly from shadowgraph images without labeled data, which first adopted a self-supervised AE to learn the light intensity fields and then employed a discrete PINN to reconstruct the refractive index fields under the constraint of the shadowgraph refractive equation. The temperature and density fields were computed based on the refractive index fields. This method effectively mitigates the sensitivity to BCs inherent in traditional numerical approaches and demonstrates high accuracy across various flow scenarios, including hot air jets, thermal plumes, and alcohol burner flames.
@@ 6.4 Others
Chi et al. 350 applied PIML to discover the heat release rate (HRR) marker for premixed NH3/H2/air flames. Specifically, a physical loss term was added to infer the free parameters in the approximation equation for HRR. Several markers were identified and then tested on various operating conditions to find the most accurate combinations. Hosseini et al. 351 used the PINNs to solve the steady-state Bratu equation arising from solid biofuel combustion theory, which was a 1D nonlinear boundary value problem. Results showed that the trained PINNs were more accurate than the decomposition method and Laplace method. Yadov et al. 352 developed the PI-LSTM to predict the linear and nonlinear response of a premixed laminar flame to incoming velocity perturbations. The low-frequency limit for perfectly-premixed flames was incorporated as a physical loss term, which effectively improved the prediction accuracy of the flame transfer functions in the minimal data limit. Kim et al. 353 developed the PI-GNN to predict the cetane number (CN) with systematic data quality analysis. Based on the physical insights, the CN-related physical properties were embedded in the global state input, thus improving the model accuracy.
@@ 7 Public resources
This section will introduce some public resources that can promote the development of PIML for combustion, including the dataset/benchmarks, simulation tools, and PIML tools. We list these resources as references not only for combustion researchers but also for AI researchers.

### v3

@@ 5.5 Combustion super-resolution via discrete PINNs
As mentioned in Section 2.3.1, discrete PINNs can be applied to SR reconstruction tasks. This field-to-field modeling paradigm is particularly suited for SR reconstruction tasks, offering advantages such as enhanced scalability for high-dimensional data, improved generalization to complex working conditions, and efficient handling of structured grid-based simulation results. Some studies on discrete PINNs for SR reconstruction of laminar, turbulent, and supersonic combustion are illustrated in Figure 11.
Figure 11: Discrete PINNs for SR reconstruction of laminar, turbulent, and supersonic combustion. (a) CRF-PINN for SR reconstruction 315. (b) The architecture of PI-ESR-GAN 316. (c) 3D volumetric SR with different models 334. (d) PI-RGAN for the SR reconstruction of flow fields in RDCs 319. Figures are reprinted with permission.
Based on the multi-receptive-field CNN 335, Zhang et al. 315 developed the combustion reacting flow PINNs (CRF-PINNs) for field reconstruction of laminar and turbulent combustion. CRF-PINNs can be used in three typical scenarios of field reconstruction: cross-field translation, SR reconstruction (as shown in Figure 11(a)), and denoising, where the reacting flow equations in a cylindrical coordinate system and the species conservation serve as the physical constraints, eliminating the need for high-fidelity labeled data during training. For the laminar Bunsen flame, with one of the velocity, temperature, and species fields, CRF-PINNs can accurately reconstruct the other two. The cross-field translation was also demonstrated for unsteady Bunsen flames. For SR reconstruction and denoising, CRF-PINNs showed superior performance over the corresponding data-driven baseline models. For turbulent flames, the turbulent viscosity and the turbulence-chemistry interaction coefficient are learned by the NN along with unknown high-fidelity fields. The Sandia flame F 329 was tested for sparse reconstruction from partial, experimental measurements. The 13 measurement points along the centerline were used for evaluation, demonstrating satisfactory model accuracy. In summary, this work proves the feasibility of discrete PINNs for various scenarios of field reconstruction, and again showcases the treatment of turbulence and chemistry in PINNs.
@@
Given the inherent complexity of combustion modeling, researchers have further augmented the regular discrete PINNs with generative modeling techniques, such as GANs. Compared to discriminative modeling approaches, generative modeling demonstrates superior capability in capturing small-scale turbulence dynamics 316 and refining predictions of discontinuous features 318. Bode et al. 316 first used the physics-informed enhanced super-resolution generative adversarial network (PI-ESR-GAN) to model turbulent combustion. As shown in Figure 11(b), the PI-ESR-GAN framework combines unsupervised deep learning with physics-informed loss functions, employing a two-step training methodology to enhance network generalization, particularly for extrapolation tasks. In decaying turbulence tests involving turbulent mixing, the model demonstrated strong a priori and a posteriori performance, with the physics-informed continuity loss term proving essential. The PI-ESR-GAN methodology has been extended to diverse combustion configurations, demonstrating remarkable versatility. In spray combustion applications, the model was successfully implemented for different cases, with validation against experimental data revealing superior predictive capability compared to classical LES approaches 336. For lean premixed combustion and gas turbine combustors, by developing a modified architecture for finite-rate chemistry flow, the enhanced model demonstrated superior accuracy in both a priori and a posteriori validations while maintaining computational efficiency 337. For turbulent non-premixed flames, a multi-mesh strategy was developed to optimize prediction accuracy across different grid resolutions, enabling reliable extrapolation to untrained Reynolds numbers 338. The approach was subsequently extended to premixed combustion 317. For that, the physical information in the loss function were adjusted, the training process was smoothed, and especially the effects of density changes were considered. When applied to DNS data of turbulent premixed flame kernels, the enhanced model achieved remarkable accuracy in reproducing fine-scale dynamics on coarse computational meshes. This advancement enables more efficient statistical analysis, while significantly reducing computational costs compared to traditional approaches. Using high-fidelity DNS data, Chung et al. 334 benchmarked the scaling behaviour of five DL models on the 3D combustion SR task, as shown in Figure 11(c). One of the models used the ESR-GAN (referred to as RRDB in the paper) and incorporated a physics-based loss (i.e., physics-informed loss in a broad sense), which penalized the gradients of channel variables in a supervised manner. It was demonstrated that the benefits of this loss can persist with increasing model size.
@@
For supersonic combustion, Wang et al. 318 developed the PI-GAN for flow field SR reconstruction on RDCs. The model utilized a GAN architecture enhanced with additional physical loss terms to ensure the outputs adhere to fundamental physical principles. Compared to interpolation and other methods, PI-GAN greatly improved the resolution of the flow field while accurately capturing detailed flow structures. Subsequently, Wang et al. 319 proposed an advanced physics-informed recurrent GAN (PI-RGAN), as shown in Figure 11(d). This innovative framework incorporated temporal derivatives as intermediate variables during the data generation process within a recurrent architecture, while simultaneously employing spatial derivatives of the output to establish comprehensive physical constraints. Experimental analysis revealed some features of the proposed model. The PI-RGAN model could well capture the detailed flow field structures and temporal characteristics in RDCs. Moreover, the introduction of a recurrent structure in the PI-RGAN model enhanced the model robustness against noisy inputs.
6 PIML for other combustion scenarios
@@
Some studies on PIML for combustion cannot be classified to the previous two sections but are also important and enlightening. Therefore, they will be introduced in this section, which are categorized into four parts: fire safety, thermoacoustics, combustion diagnostics, and others. Some of these studies are illustrated in Figure 12.
Figure 12: PIML for fire safety, thermoacoustics, combustion diagnostics, and others. (a) PINNs for modeling and prediction of lithium-ion battery thermal runaway 339. (b) PINNs to learn thermoacoustic interactions in combustors 340. (c) The structure of Lambert-Beer law PINN (LB-PINN) 341. (d) PINNs reconstruct density and temperature fields from shadowgraph images of alcohol burner flames 342. (e) PIML to discover the HRR markers for premixed NH3/H2/air flames 343. Figures are reprinted with permission.
@@ 6.1 Fire safety
Fire safety constitutes a critical discipline focused on the prevention, detection, and mitigation of uncontrolled fires, aiming to protect life, property, and the environment against fire-related hazards. However, predicting their complex dynamics remains a challenge due to the multiscale and multiphysics processes, including turbulence, heat and mass transfer, soot formation, radiation, and possible explosion. Given the advantages of PIML in integrating data and physics, researchers are exploring its applicability in advancing fire science and enhancing fire safety.
@@
Bottero et al. 344 replaced the PDE solver in the fire simulator to achieve real-time prediction of wildfire spread, where the level set equation was utilized. Tests on an ideal case and a real wildfire both demonstrated that the PINNs were able to capture the shape of the firelines. Dabrowski et al. 345 used PINNs to solve the level set equation for modeling the wildfire fire-front. A forecast likelihood was introduced to deal with the issue that PINNs failed to maintain temporal continuity when there are extreme changes in exogenous forcing variables. Moreover, the B-PINN was incorporated to provide UQ in fire-front predictions. The models were successfully applied to modeling two recorded grassland fires. Vogiatzoglou et al. 346 applied PINNs to infer the parameters in the wildfire spread model, consisting of energy and mass fraction equations. The parameters included the dispersion coefficient vector, mean gaseous velocity vector, and the overall heat transfer coefficient, so the number of parameters was three in 1D cases and five in 2D cases. Five cases were tested, including both clean and noisy data, and both synthetic and real-world data, demonstrating the accuracy of parameter inference and wildfire prediction. Some other wildfire-related applications do not focus on the spread prediction. For post-fire applications, Seydi et al. 347 used PIML to estimate the soil burn severity and reveal the intricate relationships between soil and vegetation burn severities. For firefighting, the AI Firefighter 348 was proposed to predict the optimal firefighting strategy given an arbitrary terrain and an arbitrary fire distribution, where the NN combined PIML with a decision-making process.
@@
Other applications are not related to wildfire. As illustrated in Figure 12(a), Kim et al. 339 used PINNs to model and predict the lithium-ion battery thermal runaway, which is an important topic of fire safety. The PDEs of temperature and lithium concentration served as the governing equations, which was a DR system. The model applicability to semi-supervised learning, unsupervised learning, and surrogate modeling was investigated, showing that PINNs exhibited better performance than purely data-driven NNs. Lu et al. 111 used the PI-DeepONet for surrogate modeling of the radiative heat transfer, where the radiation transfer equation (RTE) and its BCs were considered as the physical losses. The model took the location and propagation direction as trunk net inputs, the parameters (such as the absorption coefficients) as branch net inputs, so the solution operator of the RTE could be approximated. The surrogate model offers significant speedup over traditional methods, thus possessing the potential to replace the traditional RTE solvers in industry-scale fire simulations, where the solution of RTE is often a major performance bottleneck. Parsa et al. 349 applied PINNs to predict the transient burning rates of non-charring materials, where the energy equations in the temperature rise phase and burning phase were considered in the loss function. Results showed that PINNs could effectively capture the intricate dynamics of transient burning behavior and were more efficient than traditional numerical methods.
@@ 6.2 Thermoacoustics
Thermoacoustics concerns the two-way coupling between unsteady heat release and acoustic pressure waves, a fundamental phenomenon in combustion engines and combustors. Understanding this interaction is critical for both harnessing its potential in efficient thermoacoustic engines and mitigating the dangerous pressure oscillations that can lead to system failure. The predictive modeling of thermoacoustic instabilities remains challenging due to its nonlinear, multiscale, and multiphysics nature, which are now being addressed by PIML.
Ozan et al. 350 used PINNs to learn the nonlinear limit cycles and adjoint limit cycles occurred in thermoacoustic oscillations. The physics was informed in two ways: (1) the governing equations that are derived from mass, momentum and energy conservation were integrated into the loss function, and (2) the periodicity was enforced using periodic activation functions in the NN. It was demonstrated that the PDE losses could promote the model performance in noisy-data and scarce-data situations, and only the NN with periodic activation functions could accurately extrapolate to the validation set. Mariappan et al. 340 used PINNs to learn thermoacoustic interactions in a bluff body anchored flame combustor, where three PDEs arising from the acoustic model and the van der Pol (VDP) oscillator model served as the physical losses. Based on the available data of the acoustic pressure fluctuations at three locations and the total flame heat release rate, the model aimed to identify four physical parameters and recover the entire fields of the acoustic pressure, velocity, and total heat release. Five cases were tested in the order of increasing complexity with the final case using the experimental data. As shown in Figure 12(b), the satisfactory results indicated the potential of PINNs to improve existing or design new thermoacoustically stable and structurally efficient combustors. Xie et al. 351 extended the VDP equation by introducing three parameters to enhance the model applicability, and then a PINN was applied to infer these parameters, where the data were generated for a horizontal Rijke tube with acoustically compact heat sources. The optimized PINN-EVDP model was shown to be capable of replicating the thermoacoustic system instability behaviour. Further investigations were conducted to compare the Hopf bifurcation and amplitude death characteristics of the coupled extended VDP system with the coupled theoretical Rijke tube systems, showing the physical significance of the extended VDP system in the field of thermoacoustics.
@@ 6.3 Combustion diagnostics
Combustion diagnostics encompasses the techniques and methods for quantitatively measuring key parameters in combustion processes and fields. Acquiring high-fidelity experimental data is indispensable for understanding fundamental combustion physics, validating numerical models, and guiding the design of clean and efficient combustion systems. However, these measurements are challenged by the harsh conditions of practical combustors, limited optical access, and the ill-posed nature of inferring quantitative data from indirect signals like flame emissions. Since PIML has been widely used to aid experimental fluid mechanics (see Section 2.4), combustion researchers have also been exploring its potential for combustion diagnostics.
Wang et al. 341 used PINNs to reconstruct the 3D soot volume fraction in flames, where the NN took the 2D images of the accumulated projection value of the extinction coefficients as input, and output the local 3D extinction coefficients, as shown in Figure 12(c). The losses included the physical loss of the Lambert-Beer law, the boundary loss that constrained the positivity of the extinction coefficients, and the smoothness loss as a regularization term. Tested on two numerical cases and an experimental case, the model showed superior performance compared to conventional reconstruction methods. Later, Wang et al. 352 used PINNs to simultaneously reconstruct the soot temperature and volume fraction fields in experimental laminar ethylene flames. The inputs of the NN were the 2D matrix of the experimental soot radiation field and flame thickness in the line-of-sight (LoS), and the outputs were soot temperature, volume fraction fields, and an auxiliary variable. The physical loss was a first-order PDE for the auxiliary variable, which was derived from the LoS soot radiation integral equation. It was demonstrated that the physical constraint improved not only the predictive capability but also the adaptability and reliability across various combustion conditions.
Wang et al. 342 proposed a physics-informed shadowgraph network to reconstruct both density and temperature fields directly from shadowgraph images without labeled data, which first adopted a self-supervised AE to learn the light intensity fields and then employed a discrete PINN to reconstruct the refractive index fields under the constraint of the shadowgraph refractive equation. The temperature and density fields were computed based on the refractive index fields. This method effectively mitigates the sensitivity to BCs inherent in traditional numerical approaches and demonstrates high accuracy across various flow scenarios, including hot air jets, thermal plumes, and alcohol burner flames, as shown in Figure 12(d).
@@ 6.4 Others
Chi et al. 343 applied PIML to discover the heat release rate (HRR) marker for premixed NH3/H2/air flames. Specifically, a physical loss term was added to infer the free parameters in the approximation equation for HRR, as illustrated in Figure 12(e). Several markers were identified and then tested on various operating conditions to find the most accurate combinations. Hosseini et al. 353 used the PINNs to solve the steady-state Bratu equation arising from solid biofuel combustion theory, which was a 1D nonlinear boundary value problem. Results showed that the trained PINNs were more accurate than the decomposition method and Laplace method. Yadov et al. 354 developed the PI-LSTM to predict the linear and nonlinear response of a premixed laminar flame to incoming velocity perturbations. The low-frequency limit for perfectly-premixed flames was incorporated as a physical loss term, which effectively improved the prediction accuracy of the flame transfer functions in the minimal data limit. Kim et al. 355 developed the PI-GNN to predict the cetane number (CN) with systematic data quality analysis. Based on the physical insights, the CN-related physical properties were embedded in the global state input, thus improving the model accuracy.
@@ 7 Public resources
This section will introduce some public resources that can promote the development of PIML for combustion, including the dataset/benchmarks, simulation tools, and PIML tools, some of which are illustrated in Figure 13. We list these resources as references not only for combustion researchers but also for AI researchers.
Figure 13: Some public resources for PIML for combustion, including the dataset/benchmarks, simulation tools, and PIML tools.